# ORBIT-2: Scaling Exascale Vision Foundation Models for Weather and Climate Downscaling

### v1

@@ III-B TILES: Tilewise Efficient Sequence Scaling Algorithm
While Reslim significantly reduces computation by operating on low-resolution and compressed inputs, it does not resolve the inherent quadratic complexity of self-attention. As resolution increases, this limitation becomes a bottleneck. To address this, we introduce the Tilewise Efficient Sequence Scaling Algorithm (TILES), a scalable sequence processing strategy that reduces attention complexity from quadratic to linear and enables efficient parallelization across GPUs.
@@
TILES is motivated by the spatial locality property of downscaling, where the downscaling for each high-resolution pixel is primarily influenced by spatially nearby coarse-resolution inputs. This “point spread” effect, well studied in the remote sensing literature [37, 38, 39], implies that long-range pixel correlations can be safely ignored without affecting downscaling accuracy. For example, neighboring pixels such as the green and yellow ones in Fig. 4(a) have high mutual influence, while distant pairs (e.g., green and blue) contribute minimally or no influence to each other’s predictions.
@@
Leveraging this locality, TILES partitions each input and downscaling output into spatial tiles, assigning each tile to a separate GPU. Each GPU then performs downscaling separately for its assigned tile, and self-attention is restricted within each tile, preserving local context while ignoring long-range dependencies across tiles. This tilewise downscaling reduces self-attention complexity from quadratic to linear. More specifically, the computation complexity is $O(\frac{N^{2}}{T})$, where $N$ is the number of image patch tokens and $T$ is the number of tiles. For fixed-size tiles, $T$ increases proportionally with $N$, making the overall complexity linear.
@@
However, strict tiling introduces border artifacts, as pixels near the border of each tile lack context from neighboring tiles. To mitigate this, TILES introduces halo padding, where each tile is padded with a fixed-width halo (shown in gray in Fig.4(b)) that overlaps adjacent tiles. For instance, the red region in Fig.4(a) is shared between Tile 1’s halo and Tile 2’s core region both shaded in red in Fig.4(b). This overlap restores spatial continuity across tile boundaries, ensuring that border pixels—such as the yellow pixel—receive complete local context. Note that the halo width is determined empirically. Larger halos improve accuracy but increase computation, while smaller halos reduce cost but risk accuracy loss.
@@
After each GPU independently downsamples its tile, the halo regions are discarded, and the non-padded tile outputs are stitched together to form the final high-resolution output. Since each GPU processes a different tile, leading to different gradient and model parameter update, gradients from all GPUs are averaged to maintain the model consistency across GPUs and global optimization. This inter-GPU communications, however, have minimal communication frequency and overhead as it takes place only once per data batch.
@@
Figure 5: Orthogonal levels of parallelisms mapped to supercomputer at cluster, node, and device levels.
@@
Dataset name Region Resolution (km)
Input
Vars
Output
Vars
Sample Size (in/out)
# Sample
Pairs
Size
@@
TABLE I: Datasets used for pretraining, fine-tuning, and inference. Each entry specifies the downscaling resolution, input/output variable counts, sample dimensions, number of training samples, and total storage size.
III-C Orthogonal Parallelisms
TILES efficiently scales sequence length, enabling high-resolution downscaling. However, it does not address model size scaling. To support both large foundation models and global high-resolution downscaling, TILES must be integrated with complementary model-parallel strategies. Since TILES and model parallelism target orthogonal goals, with TILES for sequence length and model parallelisms for model size, they can be combined seamlessly. This results in a unified framework incorporating four distinct parallelism strategies:
  • Fully Sharded Data Parallelism (FSDP) [31]: shards both data and model parameters across GPUs, but requires temporarily gathering the full model during forward and backward passes. Requires moderate communication overhead.
  • Tensor Model Parallelism [32]: Only shards model parameters and keep parameters sharded throughout training. Requires most communication overhead.
  • Distributed Data Parallelism (DDP) [40]: distributes only the training data without sharding model parameters. Requires least communication overhead.
Once all parallelisms are established, each GPU receives a subset of the model and data that can be optimized with Flash Attention [41] to reduce memory use and cache misses, as detailed in the next subsection. Within each GPU, streaming multiprocessors (SMs) are mapped to Flash Attention cache blocks, executing vector operations in parallel within each block. Meanwhile, CPUs asynchronously load data and construct quad-trees to track the spatial layout of adaptively compressed patches described in Fig. 3.
III-D Optimizations
To further boost performance, we applied these optimizations:
Hybrid-OP Parallelism. We adopt the Hybrid-OP optimization technique from ORBIT [17], which leverages the mathematical structure of matrix chain multiplication to shard model parameters in alternating row and column dimensions. This optimization combines tensor model parallelism with FSDP, achieving superior scalability with reduced communication overhead and frequency compared to without Hybrid-OP.
Flash Attention. To accelerate self-attention computation, we use Flash Attention [41], which applies a cache-blocking technique to minimize memory access to GPU global memory. By maximizing data reuse from high-bandwidth on-chip cache, Flash Attention significantly improves compute throughput through higher cache hit rates and faster memory access.
Mixed Precision and Layer Wrapping. We further utilize BFLOAT16 mixed-precision to speed up training while reducing memory usage. To address numerical instability—where gradients with extreme magnitudes may underflow or overflow in BFLOAT16—we apply PyTorch’s dynamic gradient scaling [42]. This technique automatically rescales gradients into a representable range and reverses the scaling during parameter updates, ensuring numerical stability.
To further reduce communication cost, we apply FSDP in a layer-wise fashion [17]. Instead of sharding all model layers in a single instance, parameters are sharded one layer at a time. This reduces synchronization overhead and memory use.
(a) Reslim Architecture Speedup Comparison with ViT
ViT 9.5M 622$\rightarrow$156 24,576 1 7.3e-4 1 35.0 0.94
Reslim 9.5M 622$\rightarrow$156 24,576 1 1.1e-6 660 36.7 0.96
ViT 9.5M 112$\rightarrow$28 777,660 1x 1 OOM NA NA NA
Reslim 9.5M 112$\rightarrow$28 777,660 1 1.2e-3 NA 37.6 0.96
Reslim 9.5M 112$\rightarrow$28 777,660 1 3.6e-4 3.3 37.7 0.96
Reslim 9.5M 112$\rightarrow$28 777,660 16× 1 1.8e-4 6.6 37.8 0.96
Reslim 9.5M 112$\rightarrow$28 777,660 32× 1 1.7e-4 7.1 37.9 0.96
Reslim 9.5M 112$\rightarrow$28 777,660 4 8.0e-4 1.5 37.7 0.96
Reslim 9.5M 112$\rightarrow$28 777,660 16 6.3e-4 1.9 37.7 0.96
Reslim 9.5M 112$\rightarrow$28 777,660 36 7.4e-4 1.6 37.7 0.96
TABLE II: (a) Computing performance comparison between ViT and Reslim at 128 GPUs. For 622$\rightarrow$156 km downscaling, Reslim achieves a 660× speedup over conventional ViT while maintaining accuracy. (b) Illustrates Reslim’s performance gains at varying adaptive compression rates and tile counts, relative to a Reslim baseline without compression or tiling.
@@ III-D Optimizations
TABLE III: Maximum sequence length scaling across architectures, model sizes, compression, tiles and GPU count.
IV How Performance Was Measured
Model Configuration. All experiments in Sec. V use four model configurations: 9.5M (256-dim embedding, 6 layers, 4 heads), 126M (1024-dim, 8 layers, 16 heads), 1B (3072-dim, 8 layers, 24 heads), and 10B (8192-dim, 11 layers, 32 heads) parameters.
System Details. Experiments were conducted on the Frontier supercomputer at Oak Ridge National Laboratory. Each node consists of one 64-core AMD EPYC CPU and 8 GPUs (64 GB each), organized into 4 MI250X cards with two GPUs per card. GPUs on the same card communicate via Infinity Fabric CPU-GPU, while all four MI250X cards are connected via 50 GB/s GPU-GPU Infinity Fabric. Nodes are interconnected using 100 GB/s Slingshot-11. The software stack includes PyTorch v2.6, ROCm v6.2.4, and libfabric v1.22.
[…]
Fine-tuning is performed over the US by using both ERA5 and DAYMET at 28 km as inputs and 7 km DAYMET as output ground truth. The fine-tuning dataset is split into training, validation, and testing in the same way as pretraining.
After fine-tuning, we apply the trained model for inference to downscale global ERA5 precipitation data from 28 km to 7 km resolution, and evaluate against IMERG observation dataset [46], a NASA product for global precipitation observation. We compare the model’s daily aggregated 7 km predictions with IMERG globally. Since inference does not include bias correction—and both ERA5 and IMERG contain uncertainties—perfect alignment is not expected.
Performance Metrics. The total number of floating point operations (FLOPs) of the systems was collected via the Microsoft Deepspeed Profiler [47] and we only gathered the FLOPs on GPUs. Only the mixed-precision BFLOAT16 results were reported. We measured following performance:
@@ IV How Performance Was Measured
Dataset, Source Code and Model Checkpoint Release. We will publicly release the datasets, source code and trained model checkpoints soon and update this manuscript.
V Performance results
Figure 6: (a) TILES sequence scaling algorithm speedup across GPUs, compared to an 8-GPU baseline that does not utilize tiling. (b) Strong scaling efficiencies up to 4096 nodes (32,768 GPUs) for various model sizes, maintaining a strong scaling efficiencies of 92-98% at 4096 nodes.
V-A Speedup Ablation Studies
@@
(a) Evaluation metrics for minimum temperature (Kelvin)
Model Size R2 RMSE RMSE $\sigma_{1}>68\%$ RMSE $\sigma_{2}>95\%$ RMSE $\sigma_{3}>99.7\%$ SSIM PSNR
9.5M 0.991 3.812 4.652 9.704 15.497 0.958 29.02
@@
(b) Evaluation metrics for total precipitation (millimeter/day)
9.5M 0.975 0.146 0.166 0.344 0.449 0.931 29.03
@@
TABLE IV: Comparison of downscaling accuracy for temperature and precipitation over the U.S. using models with 9.5M and 126M parameters. Results highlight performance gains from increased model capacity.
The seventh column of Table II(a) reports the average time to downscale each hourly sample. The eighth column shows the speedup from Reslim relative to the ViT baseline. Notably, the Reslim architecture avoids expensive upsampling operations by operating directly on low-resolution inputs, resulting in significant computational savings. For the smaller 622$\rightarrow$156 km task, Reslim achieves a 660× speedup over ViT at the same number of GPUs while maintaining comparable accuracy, as measured by PSNR and SSIM. This demonstrates the effectiveness of Reslim’s residual learning design and Bayesian training loss in maintaining predictive accuracy while reducing computational cost. For the larger 112$\rightarrow$28 km resolution task, the ViT model fails due to out-of-memory (OOM) errors. Consequently, a direct speedup comparison is not available, although Reslim completes the task efficiently and maintains high accuracy.
Table II(b) explores further speedup gains from adaptive spatial compression and sequence tiling, compared to the Reslim baseline (Table II(a), row 5), all using 128 GPUs. Adaptive compression with a 32× sequence length reduction yields up to a 7.1× speedup with no loss in PSNR or SSIM. Further compression yields diminishing returns due to quad-tree overhead. Tiling provides up to 1.9× with 16 tiles per sample. Further tiling introduces excessive halo padding overhead and degrades computing performance. Accuracy remains stable across all settings.
V-B Maximal Sequence Length Scaling
Table III presents sequence length and resolution scaling performance of various model architectures and strategies, demonstrating how the combination of spatial compression, tiling, and the Reslim architecture enables extreme sequence lengths. We achieve sequence lengths of up to 4.2 billion tokens (global downscaling resolution of 0.9 km) for a 9.5M parameter model and up to 671 million tokens (global resolution 2.3 km) for a 10B parameter model. These results surpass the state-of-the-art in sequence scaling by more than 22,000×, compared to state-of-the-art sequence parallelism of 188K tokens [22], and the Swin Transformer at 147K tokens [27].
All experiments utilize 23 input variables (12 atmospheric, 6 surface, and 5 static) and produce 18 output variables (excluding static inputs). Using a standard ViT with 9.5M parameters, the maximum sequence length is limited to 25K tokens (coarse 156 km global resolution) when using 8 GPUs. Scaling this ViT model to 10B parameters results in an out-of-memory (OOM) error, making global downscaling infeasible.
@@ V-B Maximal Sequence Length Scaling
(a) Power spectrum of minimum temperature downscaling
(b) Daily total precipitation vs. observation
Figure 7: (a) Power spectra of downscaled minimum temperature using 9.5M and 126M parameter models, showing improved high-frequency fidelity with increased model capacity. (b) Daily total precipitation over the U.S. on January 1, 2020: (Left) 7 km DAYMET ground truth, and (Right) 7 km ORBIT-2 downscaling using the 126M model. ORBIT-2 accurately reconstructs fine-scale precipitation structures.
When combining Reslim with both spatial tiling (16 tiles per sample) and adaptive spatial compression (4x) techniques, substantial improvements are obtained. With these methods, the model achieves 1.1B tokens on only 8 GPUs, corresponding to downscaled output of size [11520, 23040, 18]. This result is made possible through several key compression techniques:
@@ V-C TILES Sequence Scaling Speedup
V-D Strong Scaling Efficiencies & Throughput
Fig. 6(b) presents strong scaling performance across model sizes (9.5M to 10B parameters) using the same dataset as in Fig. 6(a). Experiments were conducted at scales of 64, 256, 1024 and 4096 nodes with 8 GPUs for each node, employing all four forms of orthogonal parallelism (see Sec. III). Each point in the figure reports the average runtime in second per hourly sample with a data label for corresponding strong scaling efficiency, relative to each model’s baseline performance at 64 nodes (512 GPUs).
All model sizes maintain high strong scaling efficiencies between 92–98%. The smallest model (9.5M) underutilizes hardware at large scales due to insufficient computing with small model, with 2.5e-6 seconds per sample and a sustained computing throughput at 363 PetaFLOPS at 4096 nodes (32,768 GPUs). In contrast, larger models saturate compute resources: the 126M, 1B, and 10B models reach sustained throughputs of 1.3, 1.5, and 1.8 ExaFLOPS, respectively, at 4096 nodes. These results demonstrate the strong scalability of Reslim and the effectiveness of orthogonal parallelism for exascale climate downscaling.
@@ V-D Strong Scaling Efficiencies & Throughput
[controls=play, stop, buttonsize=1.0em, autoplay, loop, width=0.95]10figs/precipitation_animation/frame_0029
Figure 8: Playable animation comparing daily total precipitation at global scale for July 2020. (Left) Coarse-resolution ERA5 28km input; (Middle) IMERG 7km ground truth; (Right) ORBIT-2 downscaled prediction at 7 km. The animation demonstrates ORBIT-2’s ability to recover fine-scale features from coarse inputs. For optimal playback, please view in Adobe Acrobat.
V-E Fine-Tuning & Inference

### v2

@@ 4.2. TILES: Tilewise Efficient Sequence Scaling Algorithm
While Reslim significantly reduces computation by operating on low-resolution and compressed inputs, it does not resolve the inherent quadratic complexity of self-attention. As resolution increases, this limitation becomes a bottleneck. To address this, we introduce the Tilewise Efficient Sequence Scaling Algorithm (TILES), a scalable sequence processing strategy that reduces attention complexity from quadratic to linear and enables efficient parallelization.
@@
TILES is motivated by the spatial locality property of downscaling, where the downscaling for each high-resolution pixel is primarily influenced by spatially nearby coarse-resolution inputs. This “point spread” effect, well studied in the remote sensing literature (Huang et al., 2002; Wang et al., 2020a, b), shows that the correlation between pixel pairs decay rapidly when the spatial distance. Consequently, long-range pixel interactions that have weak point spread effect can be safely ignored without affecting downscaling accuracy. For example, neighboring green and yellow pixels in Fig. 4(a) have high mutual influence, while distant pairs (e.g., green and blue) contribute minimally or no influence to each other’s predictions.
@@
Leveraging this property, TILES partitions both inputs and downscaled outputs into spatial tiles, assigning each tile to a separate GPU. Each GPU then performs downscaling separately for its assigned tile, and self-attention is restricted within each tile, preserving local context while ignoring long-range dependencies across tiles. This tilewise downscaling reduces self-attention complexity from quadratic to linear. More specifically, the computation complexity is $O(\frac{N^{2}}{T})$, where $N$ is the number of image patch tokens and $T$ is the number of tiles. For fixed-size tiles, $T$ increases proportionally with $N$, making the overall complexity linear.
@@
However, strict tiling introduces border artifacts, as pixels near the border of each tile lack context from neighboring tiles. To mitigate this, TILES introduces halo padding, where each tile is extended with a fixed-width halo (shown in gray in Fig.4(b)) that overlaps adjacent tiles. For instance, the red boundary region in Fig.4(a) between Tiles 1 and 2 is duplicated, with one copy in Tile 1’s halo and another copy in Tile 2’s non-halo interior, both shaded in red in Fig.4(b). This duplication restores spatial continuity across tile boundaries, ensuring that border pixels—such as the yellow pixel—receive complete local context. Note that the halo width is tuned empirically. While in principle it can be derived directly from the data sensor’s point spread function, in practice we begin with minimal halos and increase their size iteratively until both boundary artifacts and training loss are minimized.
@@
After each GPU independently downsamples its tile, the halo regions are discarded, and the non-padded tile outputs are stitched together to form the final high-resolution output. Since each GPU processes a different tile, leading to different gradient and model parameter update, gradients from all GPUs are averaged to maintain the model consistency across GPUs and global optimization. This inter-GPU communications, however, have minimal communication overhead as it takes place only once per data batch.
@@
Figure 5. Orthogonal levels of parallelisms mapped to supercomputer at cluster, node, and device levels.
Dataset Type Interval Region Res. (km) In/Out Vars Sample Dim. (in→out) # Pairs Size (GB)
@@
ERA5 $\to$ ERA5 Reanalysis Hourly Global 622→156 23 / 3 [32,64,23]→[128,256,3] 367,920 200
ERA5 $\to$ ERA5 Reanalysis Hourly Global 112→28 23 / 3 [180,360,23]→[720,1440,3] 367,920 6,328
PRISM $\to$ PRISM Observation Daily U.S. 16→4 7 / 3 [180,360,7]→[720,1440,3] 14,235 189
DAYMET $\to$ DAYMET Observation Daily U.S. 16→4 7 / 3 [180,360,7]→[720,1440,3] 14,946 200
U.S. Regional Fine-Tuning for Tmin, Tmax and Precipitation
[ERA5, DAYMET]
$\to$ DAYMET
Reanalysis
& Observation
Daily U.S. 28→7 23 / 3 [180,360,23]→[720,1440,3] 14,946 254
Global Precipitation Fine-Tuning
[ERA5, IMERG]
$\to$ IMERG
Reanalysis
& Observation
Daily Global 28→7 23 / 1 [720,1440,23]→[2880,5760,3] 8,401 745
Table 1. Datasets used for pretraining, fine-tuning, and inference. Each entry specifies the downscaling resolution, input/output variable counts, sample dimensions, number of training samples, and total storage size.
4.3. Orthogonal Parallelisms
TILES efficiently scales sequence length, enabling high-resolution downscaling. However, it does not address model size scaling. To support both large foundation models and high-resolution downscaling, TILES must be integrated with complementary model-parallel strategies. Since TILES and model parallelism target orthogonal goals, with TILES for sequence length and model parallelisms for model size, they can be combined seamlessly. This results in a unified framework with four distinct parallelism strategies:
  • Fully Sharded Data Parallelism (FSDP) (Zhao et al., 2023): shards both data and model parameters across GPUs, but requires temporarily gathering the full model during forward and backward passes. Requires moderate communication overhead.
  • Tensor Model Parallelism (Shoeybi et al., 2020): Only shards model parameters and keep parameters sharded throughout training. Requires most communication overhead.
  • Distributed Data Parallelism (DDP) (Li et al., 2020): distributes only the training data without sharding model parameters. Requires least communication overhead.
Once all parallelisms are established, each GPU receives a subset of the model and data that can be optimized with Flash Attention (Dao, 2023) to reduce memory use and cache misses, as detailed in the next subsection. Within each GPU, streaming multiprocessors (SMs) are mapped to Flash Attention cache blocks, executing vector operations in parallel within each block. Meanwhile, CPUs asynchronously load data and construct quad-trees to track the spatial layout of adaptively compressed patches described in Fig. 3.
4.4. Optimizations
To further boost performance, we applied these optimizations:
Hybrid-OP Parallelism. We adopt the Hybrid-OP optimization technique from ORBIT (Wang et al., 2024), which leverages the mathematical structure of matrix chain multiplication to shard model parameters in alternating row and column dimensions. This optimization combines tensor model parallelism with FSDP, achieving superior scalability with reduced communication overhead and frequency compared to without Hybrid-OP.
Flash Attention. To accelerate self-attention computation, we use Flash Attention (Dao, 2023), which applies a cache-blocking technique to minimize memory access to GPU global memory. By maximizing data reuse from high-bandwidth on-chip cache, Flash Attention significantly improves compute throughput through higher cache hit rates and faster memory access.
Mixed Precision and Layer Wrapping. We further utilize BFLOAT16 mixed-precision to speed up training while reducing memory usage. To address numerical instability—where gradients with extreme magnitudes may underflow or overflow in BFLOAT16—we apply PyTorch’s dynamic gradient scaling (Pytorch, 2024). This technique automatically rescales gradients into a representable range and reverses the scaling during parameter updates, ensuring numerical stability.
To further reduce communication cost, we apply FSDP in a layer-wise fashion (Wang et al., 2024). Instead of sharding all model layers in a single instance, parameters are sharded one layer at a time. This reduces synchronization overhead and memory use.
(a) Reslim Architecture Speedup Comparison with ViT
@@ 4.4. Optimizations
[…]
Figure 6. Illustration of ERA5 surface temperature downscaling: (a) input at 622 km resolution, (b) ORBIT-2’s downscaled prediction at 156 km, and (c) ground-truth at 156 km.
5. How Performance Was Measured
Model Configuration. All experiments in Secs. 6 and 7 use four model configurations: 9.5M (256-dim embedding, 6 layers, 4 heads), 126M (1024-dim, 8 layers, 16 heads), 1B (3072-dim, 8 layers, 24 heads), and 10B (8192-dim, 11 layers, 32 heads) parameters.
System Details. All experiments were conducted on the Frontier supercomputer from Oak Ridge National Lab. Each node consists of one 64-core AMD EPYC CPU and 8 GPUs (64 GB memory each), organized into 4 MI250X cards with two GPUs per card. GPUs on the same card communicate via Infinity Fabric CPU-GPU, while all four MI250X cards are connected via 50 GB/s GPU-GPU Infinity Fabric. Nodes are interconnected using 100 GB/s Slingshot-11. Software stack includes PyTorch v2.7, ROCm v6.3.1, and libfabric v1.22.
The only exception is Table 5, which compares sustained throughput on Frontier and Alps supercomputer. Alps is built on NVIDIA GH200 Grace Hopper Superchips. Each node features 4 Hopper GPUs (96 GB memory each), with every GPU paired to a 72-core Grace CPU via a 900 GB/s NVLink-C2C interface, enabling unified CPU–GPU memory. Nodes are interconnected using NVIDIA Quantum-2 InfiniBand.
[…]
For continental U.S. fine-tuning, inputs include both ERA5 and DAYMET at 28 km, with 7 km DAYMET as output ground truth. The fine-tuning dataset is split into training, validation, and testing in the same way as pretraining. The prediction output variables are daily minimal temperature (Tmin), maximal temperature (Tmax), and total precipitation. For global fine-tuning, we target precipitation only and inputs include both ERA5 and IMERG precipitation data at 28 km, downscaled to 7 km resolution. IMERG (Huffman et al., 2020) is a satellite-based precipitation observation product from NASA. Model predictions are daily aggregated and compared to IMERG globally.
@@ 5. How Performance Was Measured
Table 3. Maximum sequence length scaling across architectures, model sizes, compression, tiles and GPU count.
Figure 7. (a) TILES sequence scaling algorithm speedup across GPUs, compared to an 8-GPU baseline that does not utilize tiling. (b) Strong scaling efficiencies up to 8192 nodes (65,536 GPUs) for various model sizes, maintaining a strong scaling efficiencies of 74-98% at 8192 nodes.
Performance Metrics. The total number of floating point operations (FLOPs) of the systems was collected via the Microsoft Deepspeed Profiler (Microsoft, 2024) and we only gathered the FLOPs on GPUs. Only the mixed-precision BFLOAT16 results were reported. We measured following performance:
@@
Dataset, Code and Model Checkpoint Release. Datasets, source code, and model checkpoints from this work will be made publicly available through the official ORNL GitHub repository. The corresponding links will be added to the arXiv version of this paper once the release process is complete.
6. Computing Performance Results
@@
Model Size 9.6M 126M 1B 10B
Frontier (MI250x) 23 76 122 166
Alps (GH200) 72 212 251 368
Table 4. Sustained computing throughput (PFLOPs) comparison of Frontier and Alps on 2048 GPUs.
Model Size 9.5M 126M 1B 10B
Inference Time (s) 4.5e-3 2.6e-2 7.2e-2 5.5e-1
Table 5. Model inference speed on a single node (8 GPUs).
[…]
Table 2(b) explores further speedup gains from adaptive spatial compression and sequence tiling, compared to the Reslim baseline (Table 2(a), row 4), all using 128 GPUs. Adaptive compression with a 32× sequence length reduction yields up to a 7.1× speedup with no loss in PSNR or SSIM. Further compression yields diminishing returns due to quad-tree overhead. Tiling provides up to 1.9× with 16 tiles per sample. Further tiling introduces excessive halo padding overhead and degrades computing performance. Accuracy remains the same across all settings.
Maximal Sequence Length Scaling. Table 3 presents sequence length and resolution scaling performance of various model architectures and strategies, demonstrating how the combination of spatial compression, tiling, and the Reslim architecture enables extreme sequence lengths. We achieve sequence lengths of up to 4.2 billion tokens (global downscaling resolution of 0.9 km) for a 9.5M parameter model and up to 671 million tokens (global resolution 2.3 km) for a 10B parameter model. These results surpass the state-of-the-art in sequence scaling by more than 22,000×, compared to state-of-the-art sequence parallelism of 188K tokens (Tsaris et al., 2024), and the Swin Transformer at 147K tokens (Liu et al., 2022).
All experiments utilize 23 input variables (12 atmospheric, 6 surface, and 5 static) and produce 18 output variables (excluding static inputs). Using a conventional ViT with 9.5M parameters, the maximum sequence length is limited to 25K tokens (coarse 156 km global resolution) when using 8 GPUs. Scaling this ViT model to 10B parameters results in an out-of-memory (OOM) error, making global downscaling infeasible.
Fine-Tuning Task Model Size R2 RMSE RMSE $\sigma_{1}>68\%$ RMSE $\sigma_{2}>95\%$ RMSE $\sigma_{3}>99.7\%$ SSIM PSNR
U.S. Tmin (Kelvin) 9.5M 0.991 3.812 4.652 9.704 15.497 0.958 29.02
@@
U.S. Precipitation (mm/day) 9.5M 0.975 0.146 0.166 0.344 0.449 0.931 29.03
@@
Global Precipitation (mm/day) 9.5M 0.986 0.098 0.120 0.136 0.133 0.923 35.09
126M 0.986 0.099 0.122 0.137 0.138 0.913 35.04
Table 6. Comparison of downscaling accuracy for fine-tuning tasks for U.S. temperature (Tmin), U.S. precipitation, and global precipitation using models with 9.5M and 126M parameters. The table includes $R^{2}$, RMSE, RMSE at different percentiles ($\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$), SSIM, and PSNR.
@@
When combining Reslim with both spatial tiling (16 tiles per sample) and adaptive spatial compression (4×) techniques, substantial improvements are obtained. With these methods, the model achieves 1.1B tokens on only 8 GPUs, corresponding to downscaled output of size [11520, 23040, 18]. This result is made possible through several key compression techniques:
@@
Strong Scaling Efficiencies & Sustained Throughput. Fig. 7(b) presents strong scaling performance across model sizes (9.5M to 10B parameters) using the same dataset as in Fig. 7(a). Experiments were conducted at scales of 64, 256, 1024, 4096 and 8192 nodes with 8 GPUs for each node at Frontier supercomputer, employing all four forms of orthogonal parallelism (see Sec. 4). Each point in the figure reports the average runtime in second per hourly sample with a data label for corresponding strong scaling efficiency, relative to each model’s baseline performance at 64 nodes (512 GPUs).
All model sizes maintain high strong scaling efficiencies between 74–98%. The smallest model (9.5M) underutilizes hardware at large scales due to insufficient computing with small model, with 1.3e-6 seconds per sample and a sustained computing throughput at 670 PetaFLOPS at 8192 nodes (65,536 GPUs). In contrast, larger models saturate compute resources: the 126M, 1B, and 10B models reach sustained throughputs of 2.4, 3.5, and 4.1 ExaFLOPS, respectively, at 8192 nodes. These results demonstrate the strong scalability of Reslim and the effectiveness of orthogonal parallelism for exascale climate downscaling.
Cross Platform Throughput Comparison. Table 5 compares sustained computing throughput on Frontier (AMD MI250x) and Alps (NVIDIA GH200) supercomputers, using 2,048 GPUs across varying model sizes. The training dataset is the same as Fig. 7, downscaling ERA5 from 112 km to 28 km resolution. On Frontier, throughput scales from 23 PFLOPs for a 9.6M-parameter model to 166 PFLOPs for a 10B-parameter model. On Alps, throughput is higher, reaching 72 PFLOPs for the 9.6M-parameter model and 368 PFLOPs for the 10B-parameter model.
@@
7. Science Performance Results